OpenStudy (anonymous):

The cost for 1 1/2 pounds of candy at Mr. Bulky’s is $7. At this rate, how much is 6 pounds of candy?

OpenStudy (austinl):

You can set this up as a proportion. \(\dfrac{1.5}{7}=\dfrac{6}{x}\) You just solve for x! :)

OpenStudy (anonymous):

1.5 costs 6 dollars therefore , 7 costs x , cross multiply

OpenStudy (anonymous):

x * 1.5 = 7*6

OpenStudy (anonymous):

solve

OpenStudy (austinl):

1.5 lbs costs $7. We are looking for how much 6 lbs will cost.